a man answered 28 out 40 questions correctly. If 70 % correct is passing , did he pass the test?.
Answer by stanbon(75887)

a man answered 28 out 40 questions correctly. If 70 % correct is passing , did he pass the test?
----------------------------
28/40 = 7/10 = 0.70 = 70%
---
He squeaked by.
Cheers,
Stan H.